### v1

# How Good Are We? Evaluating Cell AI Foundation Models in Kidney Pathology with Human-in-the-Loop Enrichment

Abstract
Training AI foundation models has emerged as a promising large-scale learning approach for addressing real-world healthcare challenges, including digital pathology. While many of these models have been developed for tasks like disease diagnosis and tissue quantification using extensive and diverse training datasets, their readiness for deployment on some arguably simplest tasks, such as nuclei segmentation within a single organ (e.g., the kidney), remains uncertain. This paper seeks to answer this key question, “How good are we?”, by thoroughly evaluating the performance of recent cell foundation models on a curated multi-center, multi-disease, and multi-species external testing dataset. Additionally, we tackle a more challenging question, “How can we improve?”, by developing and assessing human-in-the-loop data enrichment strategies aimed at enhancing model performance while minimizing the reliance on pixel-level human annotation. To address the first question, we curated a multicenter, multidisease, and multispecies dataset consisting of 2,542 kidney whole slide images (WSIs). Three state-of-the-art (SOTA) cell foundation models—Cellpose, StarDist, and CellViT—were selected for evaluation. To tackle the second question, we explored data enrichment algorithms by distilling predictions from the different foundation models with a human-in-the-loop framework, aiming to further enhance foundation model performance with minimal human efforts. Our experimental results showed that all three foundation models improved over their baselines with model fine-tuning with enriched data. Interestingly, the baseline model with the highest F1 score does not yield the best segmentation outcomes after fine-tuning. This study establishes a benchmark for the development and deployment of cell vision foundation models tailored for real-world data applications.
keywords:
Foundation Models, Segmentation Performance, Active Learning, Ensemble Learning, Domain Adaptation, Kidney
[…]
1 Introduction
AI foundation models trained on massive, diverse datasets are widely applied across numerous fields, including healthcare [1, 2]. Their versatility enables these models to tackle a variety of downstream tasks, with one of the most prominent applications being digital pathology [3, 4, 5, 6]. Among many tasks in digital pathology, cell instance segmentation is often serving as the initial step in extracting biological signals crucial for accurate disease diagnosis and treatment planning [7, 8, 9, 10, 11, 12]. The accuracy of the cell or cell nuclei segmentation forms the foundation for various subsequent biological or medical analyses, including cell type classification [13], specific cell counting [14], and cell phenotype analysis [15]. It is also considered as the stepping stone for whole slide image (WSI) analysis in any biological and biomedical applications [16].
In recent years, several cell foundation models [17, 18, 19, 20] trained on large and diverse datasets—encompassing various cell types, imaging techniques (e.g., fluorescence, brightfield, phase contrast), and experimental conditions—have demonstrated promising results. While many of these models have been developed for tasks such as disease diagnosis and tissue quantification, their readiness for deployment on some of the arguably simpler tasks, such as nuclei segmentation within a single organ (e.g., the kidney), remains uncertain. This leads us to pose the first question: Is nuclei segmentation on whole slide images (WSIs) within a single organ, like the kidney, a solved problem using current cell foundation models in histopathology?
Figure 1: Overall framework. The upper panel illustrates the diverse evaluation dataset consisting of 2,542 kidney WSIs. Performance: Kidney cell nuclei instance segmentation was performed using three SOTA cell foundation models: Cellpose, StarDist, and CellViT. Model performance was evaluated based on qualitative human feedback for each prediction mask. Data Enrichment: A human-in-the-loop (HITL) design integrates prediction masks from performance evaluation into the model’s continual learning process, reducing reliance on pixel-level human annotation.
To answer the preceding question, we have chosen the kidney as the evaluation organ for this study due to its diverse cell types. According to [21], the kidney comprises at least 16 specialized epithelial cell types, along with various endothelial, immune, and interstitial cells. Additionally, kidney whole slide images and their primary staining type, Periodic Acid-Schiff (PAS), are underrepresented in the training of current cell nuclei foundation models, underscoring a gap in the research field. We evaluated the performance of current cell nuclei foundation models in kidney pathology within a multi-center, multi-staining, multi-species setting. As shown in Fig. 1a, b, and c, we construct a diverse evaluation dataset that includes kidney nuclei data from 2,542 kidney whole slide images (WSIs) sourced from humans and rodents across both public and in-house datasets. To our knowledge, the scale of this study’s kidney WSIs surpasses all publicly available labeled nuclei datasets that include the kidney, as illustrated in Fig. 1a. We conducted a comparative analysis of three widely used SOTA cell nuclei foundation models—Cellpose, StarDist, and CellViT—for kidney cell nuclei instance segmentation. Model performance was evaluated by collectively rating each model’s prediction mask as “good,” “medium,” or “bad.” Specifically, “good” predictions were defined as those capturing approximately 90% of the nuclei in an image patch, while “medium” predictions captured between 50% and 90% of the nuclei. Predictions rated as “bad” captured less than 50% of the nuclei. This quantitative definition clarifies the criteria for each rating type, based on a standard established by a renal pathologist at Vanderbilt University Medical Center with over 20 years of experience. Our evaluation results indicated that a performance gap still exists in general nuclei segmentation for kidney pathology. Among the evaluated models, CellViT exhibited the best performance in segmenting kidney nuclei, achieving 63% of “good” ratings on the evaluation dataset.
Then, a natural follow-up question is: can we further enhance the performance of these trained cell foundation models? To tackle this, we performed data enrichment by ensembling predictions from three foundation models within a human-in-the-loop (HITL) design. Specifically, the “bad” image patches, manually corrected by pathologists, incorporated the domain knowledge required by the models, while the “good” predictions distilled insights from various foundation models. We assessed the model’s performance gains from three data enrichment strategies: using “good” image patches only, using “bad” image patches only, and using a combination of both. Our experimental results demonstrated that foundation models can mutually enhance their segmentation capabilities through data enrichment strategies, with StarDist achieving the highest F1 score of 0.82. The baseline model with the highest F1 score (CellViT) does not yield the best segmentation outcomes after fine-tuning. The combination of “good” and “bad” image patches proved to be the most effective strategy. Ultimately, these improved models can facilitate more efficient workflows in annotation software, such as QuPath [22], for kidney pathology.
2 Related work
2.1 Image Processing-Based Nuclei Instance Segmentation
A key challenge in nuclei instance segmentation is separating overlapping nuclei [23, 24, 25, 26, 27, 28, 29, 30]. Solutions to this problem range from traditional image processing techniques to deep learning approaches. Traditional methods rely on image processing for feature extraction, using intensity, shape, texture, and morphological patterns to identify nuclei. For instance, traditional methods such as marker-controlled watershed segmentation have been widely adopted for their effectiveness in separating overlapping or touching nuclei by utilizing predefined markers and adaptive thresholding techniques [23, 26]. However, these methods often struggle with over segmentation or under segmentation when dealing with highly heterogeneous or densely clustered nuclei. To address these limitations, region-based active contour models and shape-prior integration have been developed, which incorporate geometric shape information to improve boundary delineation and handle object occlusions [28]. Additionally, multi-pass segmentation strategies, like the Multi-Pass Fast Watershed (MPFW) method, have demonstrated superior performance by employing iterative segmentation passes and combining gradient-based and shape-based techniques, making them particularly effective for complex cytological samples such as cervical cell images [25]. However, these image processing-based methods remain heavily reliant on hand-crafted features, requiring domain expertise and being sensitive to hyperparameters [31, 32].
2.2 Deep Learning-Based Nuclei Instance Segmentation
To address the limitations of hand-crafted feature-based methods, deep learning (DL) has enabled automatic feature extraction. As outlined in [33], current DL approaches for nuclei instance segmentation can be broadly categorized into two-stage and one-stage approaches. Two-stage methods, such as Mask-RCNN [34], first detect and localize cell nuclei using bounding boxes before refining the segmentation in a subsequent stage. Similarly, [35] uses the first stage to generate nuclei proposals and the second stage to refine nuclei boundaries. However, two-stage methods often suffer from non-standardized training and high computational costs. Additionally, overlapping nuclei can lead to incorrect segmentation being passed from the first to the second stage, requiring further postprocessing.
In contrast, one-stage methods include DL network and postprocessing into an end to end training and inference framework. For example, Micro-Net [36] modifies the U-Net architecture [37] by incorporating multi-resolution input images, enabling better performance for nuclei of different sizes. The Dist model [38] introduces an additional decoder branch to predict distance maps from nucleus boundaries to their center of mass, aiding watershed-based postprocessing. HoVer-Net [31], a SOTA method, leverages horizontal and vertical distance maps to detect nuclei edges through the Sobel operator. StarDist [18] and CPP-Net [32] use a polygon-based approach, which both show comparable results with HoVer-Net. Cellpose [17] incorporates a U-Net architecture combined with gradient flow tracking to create an end-to-end model for nuclei instance segmentation. Similarly, network architecture in CellViT [19] uses a U-shaped vision transformer (ViT) [39] combined with HoVer-Net postprocessing. Overall, one-stage models offer a balance between accuracy and efficiency, making them more suitable for practical deployment where computational resources are limited. Leveraging the ease of standardization and training offered by one-stage methods, this work also incorporates foundation models from this branch.
2.3 Human-in-the-loop
Another key factor in improving nuclei instance segmentation is data labeling, as training instance segmentation models require pixel-level annotations for every object in an image [20]. However, creating these annotations can be expensive ($10^{-2}\text{ to }10^{0}$ USD/label) [9, 40, 41], which highlights the need to reduce the marginal cost of labeling. A recent improvement in labeling involves a multi-phase human-in-the-loop (HITL) process, where pathologists refine model errors to train foundation models sequentially [42, 9, 41]. Once the model achieves human-level performance, it is used to fully automate annotation of new images without further human correction. For example, [41] uses a two-phase HITL strategy to sequentially train models for cell segmentation and tracking by alternating between model predictions and human corrections. [9] uses a three-phase HITL strategy to build a large training dataset through crowdsourced correction and expert validation. Cellpose 2.0  [42] implements a flexible HITL pipeline for fine-tuning pretrained models with minimal new annotations. This process reduces the annotation load to 100–200 ROIs (region of interest) while still achieving high-quality segmentations. However, these approaches do not leverage multiple foundation models with HITL. Motivated by this limitation, in our work, we utilized multiple foundation models with a HITL design to scale up the training dataset and further enhance the performance of the foundation models.
Figure 2: Human-in-the-loop (HITL) Data Enrichment Design. The upper panel shows the inference and curation of prediction masks from three foundation models (Cellpose, StarDist, and CellViT) in this study. First, kidney nuclei instance segmentation was performed on the evaluation dataset using three cell foundation models. Model performance was evaluated by rating each prediction mask as “good,” “medium,” or “bad” according to criteria from a renal pathologist. “Good” predictions captured approximately 90% of the nuclei in a patch, “bad” predictions captured less than 50%, and the rest were classified as “medium.” We used this rating system to both qualitatively and quantitatively evaluate and categorize each model’s predictions within our dataset. Each rating reflects the segmentation quality of the predicted mask for an image patch. Based on the results of the image patch curation, the lower panel illustrates the data enrichment strategy that utilizes these curation outcomes to enhance model performance through continuous fine-tuning. Specifically, to enrich the training dataset while minimizing pixel-level annotation, we ensembled “good” prediction masks from multiple foundation models (curation cost: 10 seconds per image) and actively corrected strongly agreed-upon “bad” samples (annotation cost: 20 minutes per image).
3 Method
This section details the performance evaluation and HITL data enrichment framework (as shown in Fig. 2) used in this study. First, we describe the construction of a diverse, large-scale dataset for evaluating foundation models. Then, we present the HITL design that utilizes evaluations of multiple foundation models to enrich the training dataset. Lastly, the continuous fine-tuning of foundation models using the enriched dataset.
3.1 Curate a Diverse Large-Scale Dataset
To provide a comprehensive assessment of their performance in segmenting kidney nuclei, we constructed a diverse evaluation dataset. Our dataset comprises both public and private kidney data, with a total of 2,542 whole-slide images (WSIs). As illustrated in Fig.1b, 57% (1,449 WSIs) come from publicly available sources, including the Kidney Precision Medicine Project (KPMP)[43], NEPTUNE [44], and HUBMAP [45, 46], while the remaining 43% (1,093 WSIs) originate from an internal collection at Vanderbilt University Medical Center. To increase the diversity of our dataset, we incorporated WSIs from both human and rodent samples. As depicted in Fig. 1c, these WSIs were stained using Hematoxylin and Eosin (H&E), Periodic acid-Schiff methenamine (PASM), and Periodic acid-Schiff (PAS), with PAS being widely used in kidney pathology but less frequently in other organs. We randomly extracted four 512$\times$512-pixel image patches from each kidney WSI at 40$\times$ magnification. Patches that were contaminated, of low imaging quality, from dead tissue, or with incorrect staining methods were discarded. This process resulted in an evaluation dataset of 8,818 image patches.
3.2 Data Enrichment with Multiple Foundation Models and HITL
As shown in Fig. 2a, nuclei instance segmentation was performed on these kidney nuclei image patches using three cell foundation models: Cellpose [17], StarDist [18], and CellViT [19]. Then, instead of directly correcting their predicted instance masks, we use a rating-based system to evaluate the predictions from the three foundation models (as shown in Fig. 2b). This curation process involves two rounds of input from human experts. In the first round of image patch curation, two pathologist-trained students evaluated the predicted masks, categorizing them as “good,” “medium,” or “bad” based on a rating standard set by a renal pathologist with 20 years of experience. “Good” predictions captured approximately 90% of the nuclei in a patch, “bad” predictions captured less than 50%, and the rest were classified as “medium.” In the second round, the experienced renal pathologist manually reviewed and validated the uncertain samples from the first round. We used this rating system to both qualitatively and quantitatively evaluate each model’s predictions within our dataset. Each rating reflects the segmentation quality of the predicted mask for an image patch. In this study, employing multiple foundation models reduces bias in image patch curation by having each patch evaluated by three models trained with different architectures and settings. This approach offers diverse perspectives, helping to identify image patches that highlight gaps between general and specific (e.g., kidney) domain performance. As a result, curated predictions from these models can be used in the next stage of the HITL process to further refine their performance in segmenting nuclei in kidney pathology.
In this work, we leveraged the image patch curation results to reduce the pixel-level labeling effort in the HITL design. Specifically, we directly utilized the “good” samples (shown as Fig. 2d) for the next round of model refinement. Including these samples and their predictions as image and pseudo-label pairs enhances the training dataset at scale. Additionally, for each foundation model, the “good” samples reflect its strength and pretrained knowledge e(curation cost: 10 seconds per image patch). Pathologists manually corrected or annotated a small set of these “bad” image patches, which can be used for the next round of training for the foundation models (annotation cost: 10 seconds per image patch).
3.3 Data-Enriched Foundation Models Fine-Tuning
Benefiting from the data enrichment provided by multiple foundation models’ curation, the models can be fine-tuned using either “good” image patches (foundation model-generated pseudo labels), “bad” image patches (pathologist-corrected), or both, to enhance kidney nuclei segmentation performance. In this section, we first introduce the three foundation models used for image patch curation and continuous fine-tuning. A key consideration in fine-tuning, particularly when using both “good” and “bad” image patches, is the imbalance between these two classes. To address this, we applied a customized weighted oversampling method.
3.3.1 Cell Foundation Models
This section outlines the three cell foundation models used in this study. Details for each foundation model can be found in Cellpose [17], StarDist [18], and CellViT [19].
  1. 1.
    Cellpose: Cellpose [17] is a generalist segmentation model that utilizes a U-Net backbone to predict the horizontal and vertical gradients of topological maps, as well as a binary map. It was trained on a diverse dataset containing over 70,000 segmented objects from various microscopy modalities, including fluorescence, brightfield, and other specialized types. The training dataset primarily consists of fluorescently labeled cytoplasm, with or without an additional nuclear channel. Additionally, 30 H&E images from MoNuSeg [47] were included in the nucleus dataset. These H&E-stained histology images contain many small nuclei, and their polarity was inverted to make them resemble fluorescence images more closely. This variety of image sources, covering cells with diverse morphologies and modalities, allowed Cellpose to generalize well across a broad range of cell types without requiring retraining or parameter adjustments. To segment individual cells, Cellpose employs a gradient tracking algorithm that clusters pixels based on their convergence to a common center, enabling precise delineation of cellular boundaries even in challenging cases with dense packing or complex morphologies.
  2. 2.
    StarDist: StarDist [18] employs a unique star-convex polygon representation for nuclei segmentation, which is particularly effective for roundish objects such as cell nuclei. The model is built on a U-Net backbone and predicts object probabilities and radial distances from the center of each nucleus to its boundary, forming star-convex polygons. StarDist was originally developed for fluorescence microscopy images. To enhance its applicability to histopathology images, the model was further trained on the Lizard dataset [48], which includes 4,981 histopathology images, each of size 256 $\times$ 256 $\times$ 3, containing six different cell types (neutrophil, epithelial, lymphocyte, plasma, eosinophil, and connective tissue cells). During the post-processing stage, non-maximum suppression (NMS) is used to remove redundant polygons, ensuring that only unique instances are retained. Additionally, test-time augmentations and model ensembling were employed to further enhance segmentation performance, making StarDist a robust solution for diverse microscopy image types.
  3. 3.
    CellViT: CellViT [19] employs a U-Net shaped hierarchical encoder-decoder Vision Transformer (ViT) backbone, designed specifically for nuclei segmentation and classification in histopathology images. The encoder utilizes pre-trained weights from a ViT trained on 104 million histological images (ViT256)[49], a model that demonstrated superior performance in cancer subtyping and survival prediction tasks. CellViT is further trained on the PanNuke dataset [50], which includes 189,744 annotated nuclei across 19 tissue types, grouped into five clinically relevant classes: neoplastic, inflammatory, epithelial, dead, and connective. The images in this dataset are captured at 40$\times$ magnification with a resolution of 0.25 µm/pixel, making it a challenging benchmark due to its diversity and class imbalance. CellViT’s post-processing pipeline follows the HoVer-Net methodology [31], using gradient maps of horizontal and vertical distances for accurate boundary delineation. Additionally, the network benefits from extensive data augmentation techniques and transfer learning strategies to overcome the scarcity of annotated medical data, achieving SOTA performance on both the PanNuke and MoNuSeg datasets.
3.3.2 Class-wise Weighted Oversampling
In this work, we addressed the imbalance between human-annotated image patches and pseudo labels from foundation models’ “good” predictions (with a ratio of about 1:50) by applying oversampling. Specifically, we implemented a customized class-based weighted oversampling method, as used in [19]. For a training dataset of $N_{\text{Train}}$ samples, the sampling weight for each training sample $i$ was calculated based on its curated class (“good” or “bad” image patches), as outlined below:
$\begin{split}w_{\text{Class}}(i,\gamma_{s})=\frac{N_{\text{Train}}}{\gamma_{s}\left(\sum_{j\in[1,N_{\text{Train}}]|\ell_{j}=\ell_{i}}1\right)+(1-\gamma_{s})N_{\text{Train}}}\end{split}$ (1)
As denoted, $w_{\text{Class}}(i,\gamma_{s})$ represents the sampling weight assigned to image patch $i$ based on its curated class. Higher weights will be assigned to underrepresented classes, ensuring they are sampled more frequently during training. $\gamma_{s}\in[0,1]$ is the oversampling parameter that controls the strength of oversampling. When $\gamma_{s}=0$, there is no oversampling, and all patches are sampled uniformly. When $\gamma_{s}=1$, the maximum balancing is applied, and patches from underrepresented classes are given the highest weights. In this work, we followed the default setting of $\gamma_{s}=0.85$ in [19]. $\sum_{j\in[1,N_{\text{Train}}]|\ell_{j}=\ell_{i}}1$, this sum counts the number of patches in the training dataset that belong to the same class as patch $i$. Specifically, $\ell_{j}$ and $\ell_{i}$ represent the class labels of patches $j$ and $i$, respectively. The numerator $N_{\text{Train}}$ ensures that the weights are normalized based on the total number of training patches. The expression $\gamma_{s}\left(\sum_{j\in[1,N_{\text{Train}}]|\ell_{j}=\ell_{i}}1\right)+(1-\gamma_{s})N_{\text{Train}}$ controls the balance between classes. When $\gamma_{s}$ is large, equation (1) gives more weight to underrepresented classes (with fewer samples), while for smaller $\gamma_{s}$, equation(1) reduces the effect of class imbalance, leading to more uniform sampling. In summary, this oversampling approach ensures that human annotated samples, such as those with challenging nuclei or lower image quality, are sampled more frequently during training, which helps improve the model’s ability to generalize across all cell nuclei in the kidney.
4 Experiments
4.1 Dataset
Evaluation Dataset. As described in the previous section, the evaluation dataset in this study is highly diverse, comprising 2,542 whole-slide images (WSIs) from both public and private sources. It includes samples from both human and rodent tissue, stained with H&E, PASM, and PAS. After discarding contaminated or poor-quality patches, we randomly extracted four 512 $\times$ 512-pixel patches from each WSI at 40$\times$ magnification, resulting in an evaluation dataset of 8,818 image patches. As illustrated in Fig.1b, 57% (1,449 WSIs) were collected from publicly available sources, including the Kidney Precision Medicine Project (KPMP)[43], NEPTUNE [44], and HUBMAP [45, 46], while the remaining 43% (1,093 WSIs) were acquired from an internal collection at Vanderbilt University Medical Center. To increase the diversity of our dataset, we incorporated WSIs from both human and rodent samples.
Fine-tuning Dataset. Building on the evaluation dataset construction described above, three foundation models were used for inference and curation on each 512 $\times$ 512 image patch. The “good” prediction masks from all three foundation models, along with their corresponding images, were retrieved as image-pseudo label pairs and added to the fine-tuning dataset. Additionally, a small set of 198 images, rated as “bad” by all foundation models, were manually annotated by pathologists and included in the fine-tuning dataset. This resulted in a total of 12,005 image-label pairs. For the training and validation split, 11,807 pairs were used for training, while a diverse set of 100 pairs was reserved for validation. Lastly, an additional 185 images, sampled from all kidney WSI sources, were annotated by pathologists and designated as the hold-out testing set to evaluate the performance of the fine-tuned foundation models. Each 512 $\times$ 512 image patch contains 50 to 300 cell nuclei, providing substantial signals for effective fine-tuning.
4.2 Experimental Design
4.2.1 Evaluation of Foundation Model Performance
The segmentation of nuclei instances from the three foundation models was evaluated as “good,” “medium,” or “bad,” following a standard established by a renal pathologist with more than 20 years of experience. We applied this rating system to quantitatively assess each cell foundation model’s predictions in our evaluation dataset. The specific definitions are as follows:
  • Good: Predictions capture approximately 90% of the nuclei in an image patch.
  • Medium: Predictions capture between 50% and 90% of the nuclei in an image patch.
  • Bad: Predictions capture less than 50% of the nuclei in an image patch.
Single Model Performance Evaluation. First, we evaluated the performance of each individual foundation model by analyzing the distribution of rating assignments across our evaluation dataset. This analysis provided insights into each cell foundation model’s performance and behavior by examining the frequency of prediction classes (“good,” “medium,” “bad”) for each foundation model.
Fused Model Performance Evaluation. Building on the evaluation of each individual cell foundation model’s performance, we conducted a joint model performance evaluation. The rating assignment for each image patch was determined by the fusion of ratings from the three foundation models. Specifically, an image patch was assigned to the “good” prediction class if any model rated it as “good.” Conversely, an image patch was categorized as “bad” only if all foundation models were assigned a “bad” rating. The remaining image patches were categorized as “medium.” Similarly, the distribution of rating classes was analyzed. The results of this analysis indicate the upper bound of applying multiple foundation models to our domain task and highlight the potential for fine-tuning these models through our data enrichment strategies. Furthermore, a taxonomy of common errors made by all cell foundation models in this study can be derived from the fused “bad” image patches.
Cross-Model Performance Evaluation. In this work, we also evaluated cross-model performance on our kidney nuclei dataset by conducting an agreement analysis among the foundation models. The steps are as follows:
  1. 1.
    Agreement Matrix. We computed the percentage of agreement between each pair of models to assess how often they assigned the same rating to image patch predictions. This matrix highlights the consistency among the foundation models in their rating assignments.
  2. 2.
    Class-Specific Agreement. Image patches were categorized into three groups: All Three Models Agree, where all models assigned the same rating; Two Models Agree, where any two models assigned matching ratings; and No Agreement, where none of the models assigned the same rating to the image patch.
4.2.2 Data-Enriched Fine-Tuning with Multiple Foundation Models
To improve kidney nuclei segmentation performance, we leveraged data enrichment from multiple foundation models’ curation outcomes and fine-tuned the models under three different settings. Each foundation model was fine-tuned using either “good” image patches (foundation model-generated pseudo labels), “bad” image patches (pathologist-corrected), or both. As previously explained, fine-tuning on corrected “bad” image patches aims to narrow the knowledge gap between domains by focusing on samples where all foundation models fail. “Good” image patches enable mutual knowledge distillation across foundation models, enhancing the training of each individual model. The details of the fine-tuning experiments are shown in the Table. 1. For each fine-tuning setting, we used different scales of the training dataset (25%, 50%, 75%, and 100%) to assess the consistency of the model’s fine-tuning. Specifically, each entry in the “Incremental Training Dataset Settings” represents the size of the scaled training dataset used for fine-tuning experiments. For example, when fine-tuning with only “good” image patches derived from foundation models’ predictions, we conducted experiments using 2,952 (25%), 5,901 (50%), 8,854 (75%), and 11,807 (100%) “good” predictions. Similar experiments were performed for all three fine-tuning strategies across the three foundation models: Cellpose, StarDist, and CellViT.
Fine-tuning Experiments Data Labeling Source Incremental Training Dataset Settings
25% 50% 75% 100%
“good” patches only Models’ Predictions 2,952 5,901 8,854 11,807
“bad” patches only Pathologists’ correction 50 99 149 198
“good” + “bad” patches Combine Both 3,002 6,000 9,003 12,005
Table 1: The table summarizes the model fine-tuning experiments conducted under three settings: using “good” image patches (foundation model-generated pseudo labels), “bad” image patches (pathologist-corrected), and a combination of both. Each data-enriched fine-tuning experiment was performed at different scales of the training dataset (25%, 50%, 75%, and 100%) for the three foundation models: Cellpose, StarDist, and CellViT. The numerical values indicate the size of the scaled training dataset for each individual fine-tuning experiment.
4.3 Evaluation Metrics
To assess the instance segmentation performance of the foundation models after fine-tuning, we used Recall (equation(2)), Precision (equation(3)), and F1 score (equation(4)) as evaluation metrics.
Recall measures the ability of the model to correctly identify all nuclei instances in an image patch.
Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}=\frac{TP}{TP+FN}$ (2)
Precision measures the proportion of correctly identified nuclei instances among all the instances predicted by the model.
Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}=\frac{TP}{TP+FP}$ (3)
F1 score is calculated as the harmonic mean of precision and recall, providing a single metric that balances both by weighting their average.
$F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\times TP}{2\times TP+FP+FN}$ (4)
Moreover, the Intersection over Union (IoU) threshold is set to 0.5 to consider a detection as a True Positive (TP). $|TP|,|FP|,\text{and }|FN|$ are the number of True Positives, False Positives, and False Negatives, respectively.
  • True Positives (TP): Matched pairs of nuclei, i.e., correctly detected nuclei.
  • False Positives (FP): Unmatched predicted nuclei, i.e., predicted nuclei without matching ground truth (GT) nuclei.
  • False Negatives (FN): Unmatched GT nuclei, i.e., GT nuclei without matching predicted nuclei.
4.4 Implementation Details
Inference: For a simple implementation of model inference using each foundation model’s released pretrained weights with GPU support, our previous work [51] provides customized object-oriented python modules. Models were trained using Python 3.9 and PyTorch 2.0.1, with GPU acceleration provided by CUDA 11.7. The experiments were conducted on an NVIDIA RTX A5000 GPU with 24GB of memory.
To maintain consistency in experimental settings and enable better comparison of fine-tuned performance, we applied the same oversampling factor $\gamma_{s}=0.85$ across all experiments that used both “good” and “bad” image patches. Due to the imbalance between “good” and “bad” image patches, the number of training epochs was set to 50 for all fine-tuning experiments involving “good” image patches, while experiments using only “bad” image patches were trained for 25 epochs. The batch size was consistently set to 16 across all experiments.
Cellpose Training: For training Cellpose using our true color images, we first converted the images into the format required for training the Cellpose nucleus model. Each image was converted into a 2-band format, where the first band represents the grayscale image and the second band consists entirely of zeros. This follows the default Cellpose nucleus model inference format, where the first channel is used for nuclei segmentation. Due to different loss convergence patterns in training Cellpose, the base learning rate was set to 0.1 for fine-tuning with only “bad” image patches and 0.001 for fine-tuning that includes “good” image patches. Other experimental settings are unchanged according to  [17].
StarDist Training: The learning rate was maintained at 0.0003 (default), with all other settings unchanged as in [18]. In this work, the fine-tuned StarDist model can be converted and loaded into QuPath extension format for easy validation and labeling purposes.
CellViT Training: The base learning rate was set to 0.0003, with a scheduling factor of 0.85 to gradually reduce the learning rate during training. HoVer-Net post-processing with equal weighting on each loss objective (excluding nuclei classification and tissue classification) was applied during training. The pretrained ViT256 encoder and decoder were trained end-to-end, with no layers frozen. All other hyperparameters remained the same as in the original setting [19].
5 Results
5.1 Evaluation of Foundation Model Performance
Individual Model Performance Evaluation. First, we present the evaluation of each individual foundation model’s performance. We analyzed the distribution of human-rated predictions for each foundation model across our evaluation dataset. From the barplot in the upper panel of Fig. 3, we can observe the distribution of the rated predictions for each foundation model in this study. For all three foundation models (Cellpose, StarDist, and CellViT), at least 40% of predictions are rated as “bad” or “medium.” at least 40% of predictions are “good.” This indicates that, despite not being specifically trained on kidney pathology nuclei data, the models retain some transferable knowledge applicable to our task. However, none of the foundation models is perfect, and the 40% performance gap highlights the ongoing need for more specialized foundation models tailored specifically to kidney pathology and their challenging cell nuclei segmentation cases. Among the evaluated cell foundation models, Cellpose and StarDist have similar amounts of “good” rated predictions, while CellViT shows a notably higher number of “good” predictions of 5,609 (63%), suggesting its potentially superior performance and adaptability in kidney pathology.
Fused Model Performance Evaluation. As shown in Fig. 3, building on the evaluation of each individual cell foundation model’s performance, we conducted a fused model performance evaluation through a data enrichment strategy. The rating assignment for each image patch was determined by the fusion of ratings from the three foundation models. This evaluation aims to demonstrate the potential of applying multiple foundation models together in our domain task. As shown, after data enrichment, the fused model performance indicates an increase in the number of “good” image patches (6,001 “good” image patches, representing 68% of the dataset) and a decrease in the number of “bad” image patches (534 image patches, representing 6% of the dataset). Consequently, the “good” image patches surpass those of each foundation model, while the “medium” and “bad” image patches fall below all three models. This highlights the expanded capacity of the foundation models and supports subsequent continual learning using data-enriched prediction outcomes.
Figure 3: Distribution of Rated Predictions from Cell Foundation Models Across the Evaluation Dataset. Each row represents the foundation model’s predictions, with three values corresponding to the number of predictions rated as “good,” “medium,” and “bad,” respectively. Then, data enrichment (shown as ”Fused” Model) was performed based on the evaluation results of individual models, resulting in an increase in “good” image patches and a decrease in “bad” image patches. Lastly, we summarized a taxonomy of “bad” image patches that all foundation models failed.
Furthermore, as illustrated in the lower panel of Fig. 3, a taxonomy of common errors made by all cell foundation models is summarized from the fused set of “bad” image patches. It is evident that certain types of nuclei and tissue conditions diminish the segmentation performance of foundation models. Specifically, long and flat nuclei, nuclei with blurred boundaries, and densely distributed nuclei within glomeruli are particularly difficult to segment accurately. Additionally, slides with faint staining, those containing fatty tissues, or slides with an excessive number of red blood cells can result in a higher incidence of model failures.
Cross-Model Performance Agreement. Next, we move beyond evaluating the performance of individual foundation models and assess their collective behavior by examining the agreement and disagreement among their predictions. Fig. 4a shows the consistency between model predictions. As illustrated, no pair of models reaches over 90% agreement or complete disagreement. The highest agreement occurs between Cellpose and StarDist (0.76), while the lowest is between Cellpose and CellViT (0.67). These non-extreme values, ranging from 0.67 to 0.76, suggest that the current SOTA nuclei foundation models in digital pathology do not generalize exceptionally well to large-scale, diverse kidney datasets. Each model exhibits its own strengths, highlighting the potential for combining multiple SOTA foundation models to improve performance in downstream kidney nuclei tasks. Fig. 4b shows the percentages of image patches where all three models agree, two models agree, or no models agree for each rated prediction class (“good”, “medium”, “bad”). It is evident that “bad” samples exhibit the highest inter-model reliability, with over 70% agreement among all three models, indicating the nature of our design that prioritizes human annotation effort for consensus “bad” image patches.
Figure 4: Cross-Model Performance Agreement. (a) shows the agreement matrix between each pair of foundation models. To further assess the cross-model performance, (b) shows the percentages of image patches where all three models agree, two models agree, or no models agree, for each prediction class (“good”, “medium”, “bad”).
To sum up, by combining the performance evaluations from the results above, we found that none of the foundation models in this study is fully capable of addressing all cell nuclei segmentation challenges in the kidney. This suggests that current foundation model-based segmentation methods are still insufficient to fully solve nuclei segmentation across all organs or cell types in digital histology.
Fine-tuned Cellpose Model F1 score Precision Recall
Cellpose Baseline 0.6748 0.7006 0.6694
good 25% 0.7297 0.8313 0.6649
good 50% 0.7357 0.8286 0.6748
good 75% 0.7422 0.8447 0.6747
good 100% 0.7529 0.8349 0.6986
bad 25% 0.5353 0.8095 0.4170
bad 50% 0.6444 0.7959 0.5592
bad 75% 0.5671 0.8401 0.4461
bad 100% 0.5807 0.8121 0.4699
good 25% + bad 25% 0.7228 0.8475 0.6466
good 50% + bad 50% 0.7270 0.8481 0.6493
good 75% + bad 75% 0.7352 0.8535 0.6953
good 100% + bad 100% 0.7350 0.8512 0.6601
Fine-tuned StarDist Model F1 score Precision Recall
StarDist Baseline 0.7380 0.9158 0.6331
good 25% 0.7926 0.8989 0.7209
good 50% 0.7931 0.9084 0.7166
good 75% 0.7895 0.9063 0.7110
good 100% 0.7943 0.8991 0.7234
bad 25% 0.8086 0.8862 0.7423
bad 50% 0.8166 0.8834 0.7696
bad 75% 0.8109 0.8864 0.7601
bad 100% 0.8229 0.8847 0.7802
good 25% + bad 25% 0.7980 0.9017 0.7249
good 50% + bad 50% 0.8110 0.9015 0.7464
good 75% + bad 75% 0.8172 0.8961 0.7588
good 100% + bad 100% 0.8182 0.8901 0.7664
Fine-tuned CellViT Model F1 score Precision Recall
CellViT Baseline 0.7838 0.8233 0.7629
good 25% 0.7836 0.8216 0.7621
good 50% 0.7828 0.8296 0.7520
good 75% 0.7854 0.8284 0.7587
good 100% 0.7794 0.8270 0.7476
bad 25% 0.7796 0.7866 0.7857
bad 50% 0.7770 0.7898 0.7791
bad 75% 0.7881 0.7770 0.8137
bad 100% 0.7861 0.7809 0.8070
good 25% + bad 25% 0.7893 0.8259 0.7695
good 50% + bad 50% 0.7895 0.8276 0.7652
good 75% + bad 75% 0.7866 0.8303 0.7580
good 100% + bad 100% 0.7952 0.8302 0.7731
Table 2: Fine-tuning Results for Cellpose, StarDist, and CellViT Models. This table consists of three subtables summarizing the instance segmentation performance metrics (F1 score, Precision, and Recall) for each foundation model across three data enrichment experimental settings (using only “good” image patches, only pathologist-corrected “bad” image patches, and a combination of both). Each subtable compares the baseline performance (highlighted in blue) of the model against its fine-tuned counterparts. The highest values of each evaluation metric for the fine-tuned models are highlighted in red.
Figure 5: Comparison of Fine-Tuning results across three foundation models. Each column represents the fine-tuning of an individual foundation model, and each row compares an instance segmentation performance metric (F1 score, Precision, or Recall) across the models. The X-axis in each plot corresponds to the incremental percentage (25%, 50%, 75%, and 100%) of the training dataset used for each fine-tuning strategy, and the Y-axis shows the performance metrics.
Figure 6: Comparison of Best Fine-Tuned Models Against Baseline Performance. Each foundation model’s baseline is highlighted in pink, while our best data-enriched fine-tuned model is shown in blue. This indicates that all models enhance segmentation performance after fine-tuning. Notably, the baseline model with the highest F1 score before fine-tuning (CellViT) does not yield the best segmentation outcomes after fine-tuning. In contrast, StarDist benefits the most from the proposed data enrichment strategies, achieving the highest F1 score of 0.8229.
5.2 Data-Enriched Foundation Models Fine-Tuning
To enhance model performance while minimizing reliance on pixel-level human annotation, we further fine-tuned the foundation models using our data enrichment strategies, which ensemble the evaluation outcomes of multiple foundation models. For each foundation model, the pretrained weights served as the baseline method. First, we evaluated the baseline performance of each cell foundation model using our hold-out testing dataset. The results of assessing the baselines for Cellpose, StarDist, and CellViT are shown in blue in Table. 2. Consistent with our previous performance evaluation from human feedback, CellViT demonstrates superior performance compared to the other two models, with an F1 score of 0.7838, followed by StarDist at 0.7380, and Cellpose at 0.6748. Although StarDist achieves a high precision score of 0.9158, its lower recall of 0.6331 suggests that it misses certain amount of kidney cell nuclei. Cellpose exhibits the largest domain gap with the lowest F1 score and shows at least 10% decrease in all evaluation metrics (F1 score, Precision, and Recall) compared to CellViT.
Then, we investigated the performance of foundation models from our data-enriched fine-tuning strategies. In Fig. 5, We grouped the fine-tuning results across all three foundation models in this work. Each column represents the fine-tuning results for an individual foundation model, while each row compares an instance segmentation performance metric (F1 score, Precision, or Recall) across the models. Different colors represent distinct fine-tuning strategies or settings. The X-axis in each plot corresponds to the incremental percentage (25%, 50%, 75%, and 100%) of the training dataset used for each fine-tuning strategy, and the Y-axis shows the performance metrics. From the fine-tuning performance trends illustrated in Fig. 5, both the Cellpose and StarDist models demonstrate significant improvements in F1 score compared to their baseline models. These gains highlight the effectiveness of the data enrichment strategies employed during fine-tuning. The details of instance segmentation performance gains are provided below with the evaluation metric values referenced in Table. 2.
  1. 1.
    For the Cellpose model, fine-tuning with only “good” image patches at all training dataset scales significantly boosts performance compared to the model baseline, with the F1 score increasing from 0.6748 to 0.7529, Precision improving from 0.7006 to 0.8447, and Recall increasing from 0.6694 to 0.6986. However, fine-tuning with “bad” image patches alone results in lower performance. Although Precision reaches high at 0.8401, Recall drops significantly, with the highest recall being 0.5592, which is over 10% below the baseline recall score. The F1 scores also decrease, with the highest value being 0.6444 below baseline F1 score at 0.6748, indicating poor performance when relying solely on these “hard” samples. Fine-tuning with a combination of “good” and “bad” image patches yields stable improvements across all metrics. This shows that the presence of “good” image patches from multiple foundation models’ predictions is crucial, which compensates for the model’s inability to learn effectively from “bad” image patches alone. While the F1 scores and Precision are slightly lower compared to using only “good” patches, they still significantly outperform the baseline.
  2. 2.
    For the StarDist model, fine-tuning trend in Fig. 5 across all data enrichment fine-tuning strategies resulted in notable instance segmentation performance gains. The F1 score increased significantly from the baseline of 0.738 to 0.8229, and Recall improved from 0.6331 to all above 0.7802. While Precision saw a slight drop from 0.9158 to around 0.89, the decrease is minimal compared to the substantial improvements in F1 and Recall. This highlights the effectiveness of each data enrichment strategy in improving segmentation quality, with “bad” samples minimizing the knowledge gap through expert labeling, and “good” samples distilling knowledge from other foundation models. This demonstrates that our HITL design, based on the curation outcomes of multiple foundation models, can effectively improve model performance.
Lastly, according to both Fig. 5 and Table. 2, CellViT shows relatively consistent F1 scores across all fine-tuning strategies. Fine-tuning with only “good” data slightly increases Precision but causes a small reduction in Recall (e.g., from 0.7629 to 0.7476 with 100% “good” data, with the highest recall being 0.7621). In contrast, fine-tuning with only “bad” data improves Recall (from 0.7629 to 0.8137) but reduces Precision (from 0.8233 to 0.7898). Combining “good” and “bad” data offers the best of both data enrichment strategies, with improved F1 scores at every training scale. For instance, using 100% “good” and “bad” image patches achieves the highest F1 score (0.7952), with balanced Precision (0.8302) and Recall (0.7731), all surpassing the baseline. Although the performance gain is not as pronounced as with Cellpose or StarDist, integrating both “good” and “bad” image patches results in the most balanced overall performance, enhancing both Precision and Recall.
Summarizing the best fine-tuned model for each cell foundation model, Fig. 6 compares these models to their corresponding baselines. This indicates that all models enhance segmentation performance after fine-tuning. Notably, the baseline model with the highest F1 score before fine-tuning (CellViT) does not yield the best segmentation outcomes after fine-tuning. In contrast, StarDist benefits the most from the proposed data enrichment strategies, achieving the highest F1 score of 0.8229.
Figure 7: Impact of Incrementally Increasing Percentage (%) of High-Quality Labels on Model Performance. The dashed line represents the baseline comparison setting, using 25% pseudo labels from multiple foundation models’ “good” predictions and 25% pathologist-corrected labels, which represent the gap of the models.
5.3 Effect of Label Imbalance on Model Performance
Following the quantitative analysis above, compared to Cellpose and StarDist, which show a significant increase in performance over their baseline methods, CellViT’s overall performance gain is not as significant. We hypothesized that this can be attributed to label imbalances in the training dataset. As shown in Fig. 3, CellViT generates the majority of the pseudo labels, leading to an imbalance between these pseudo labels and pathologist-corrected labels (representing “bad” image patches), which highlights the domain gap. Additionally, there is an imbalance within the pseudo labels themselves; the proportion of pseudo labels rated as “good” by other models, while not by CellViT, is less pronounced compared to the fine-tuning of StarDist and Cellpose. This imbalance can be observed by examining the difference in the number of “good” image patches between the fused model and the individual foundation models in Fig. 3. A larger difference indicates that data enrichment distills more image patches needed by the model.
Thus, we performed an ablation study to assess the impact of adding high-quality labels, which the foundation models need in our imbalanced fine-tuning dataset. For ease of implementation, we incrementally increased the percentage of pathologist-corrected “bad” image patches in the training dataset. The baseline comparison (dash line) used 25% pseudo labels from multiple models’ “good” predictions and 25% pathologist-corrected labels. In this study, we incrementally incorporated an additional 25% of “bad” image patches into the training dataset, presenting the model performance results as blue curves in Fig. 7. Consistent with our previous analysis, we observe that the inclusion of “bad” image patches, which are often faintly stained samples, decreases the performance of Cellpose. This suggests that fine-tuning the Cellpose model requires “good” image patches to maintain the pretrained knowledge and avoid catastrophic forgetting during our continual fine-tuning. For both CellViT and StarDist, we observe a clear pattern of increased performance over the baseline as the availability of model-required high-quality labels rises. Incorporating additional “bad” image patches leads to notable increases in the F1 score. For CellViT, when the percentage of “bad” image patches in the training dataset is increased to 50%, CellViT achieves an F1 score of 0.795, which is equivalent to its peak performance when using all “good” and “bad” image patches, as shown in Table 2. With inclusion of all “bad” image patches, it reaches highest F1 score of 0.798. This improvement is attributed not only to the inclusion of high-quality labels but also to the oversampling of these image patches in the training dataset. Overall, the results demonstrate that the data enrichment strategy of utilizing both “good” and “bad” image patches together enhances the performance of StarDist and CellViT, and oversampling the model-needed data is a crucial aspect of our approach.
Figure 8: Qualitative Results of Enhanced Kidney Cell Nuclei Instance Segmentation. The predicted nuclei and ground truth are represented as overlaid green contours on the image patch, with areas of improvement highlighted by rectangles.
5.4 Visualization of Improved Model Performance
In this section, we present the qualitative visualization results of our enhanced kidney cell nuclei instance segmentation, evaluated using the best fine-tuning strategies from three foundation models. Three example image patches (A, B, and C) are presented in Fig. 8. The upper panel for each example shows the baseline performance across all models, while the lower panel displays the segmentation performance following data-enriched fine-tuning. As shown in Fig. 8(A), our fine-tuned Cellpose model improves the instance segmentation on long and flat nuclei. (B) highlights the improvement of our StarDist model in segmenting dense nuclei in a glomerulus. The last panel shows the improvement in segmenting nuclei in PAS-stained images, which are underrepresented in the current histology dataset. Most faintly stained nuclei that were missing in the original predictions within a glomerulus are detected by the fine-tuned StarDist and CellViT models. In contrast, the fine-tuned Cellpose model generates fewer false positive predictions compared to its baseline. Additionally, as illustrated in both (A) and (B), the highlighted regions in the image patches represent knowledge gaps in the target domain where the inferior model struggles with certain types of image patches, while the other models do not. Examples of our fine-tuning results demonstrate that this less effective model can benefit from the knowledge of the other models to improve its performance. Additional visualizations of the qualitative results (D)-(L) are provided in Supplementary Figs. S1, S2, and S3 (See Supplementary Information for additional data)
6 Discussion
The evaluation of Cellpose, StarDist, and CellViT reveals that a performance gap still exists between general and kidney-specific nuclei segmentation. Cell nuclei segmentation cannot be considered a fully solved problem across all organs, tissues, and imaging modalities when utilizing off-the-shelf foundation models. Our foundation model-based data enrichment strategies aim to meet the need for a kidney-level cell foundation model and to reduce labeling costs. In this process, the collective image curation from multiple foundation models is crucial, as a single model may not capture the diversity of segmentation styles in the target kidney domain. By leveraging the versatility of multiple foundation models, annotation efforts are minimized and transformed into time-efficient image rating curation, which can be further enhanced through the model’s uncertainty analysis. In the first round of HITL fine-tuning, all three foundation models demonstrated improved performance over their baselines, with StarDist achieving the highest segmentation performance, achieving an F1 score of 0.82. Both StarDist and CellViT highlight that including “good” and “bad” image patches is the most effective training strategy in our approach. Although Cellpose shows a significant increase in performance (with the F1 score rising from 0.67 to 0.75) with our fine-tuning, it demonstrates lower efficacy when only hard image samples are included. This is primarily because the Cellpose model is designed for fluorescently labeled images, whereas the other two models have been adapted for H&E-stained images. For the training of segmentation task, 512 $\times$ 512 images were converted into grayscale nuclear bands, which can lead to information loss, particularly for hard images with very light staining colors and complex tissue morphological patterns.
A limitation of this approach is that it requires more expert knowledge and image preprocessing to enhance nuclear intensity while minimizing the noisy background in kidney whole-slide images. Another aspect that requires further improvement is the imbalance of labels in the training dataset. As investigated in our ablation study, oversampling the labels that highlight the domain gap is essential. However, this approach presents a trade-off, as labeling or curating these samples from a large-scale dataset incurs additional manual costs. In our future research, we plan to employ semi- or fully-automated model selection to help mitigate this workload.
Lastly, as the models become better adapted to the specific kidney organ and the performance gap decreases among three models after the first round of fine-tuning, we anticipate generating improved curation results, including a higher number of “good” ratings, a reduction in “bad” samples, and more accurate prediction masks in the next round of HITL fine-tuning. This improvement can be further enhanced by utilizing our improved model (e.g., StarDist) in QuPath software for label correction by pathologists.
7 Conclusion
In this work, our evaluation on a comprehensive kidney dataset indicates that cell nuclei segmentation in histopathology still needs further improvement with more specific organ-targeted foundation models. Among the evaluated models, CellViT demonstrated superior performance in segmenting nuclei in kidney pathology, achieving an F1 score of 0.78 in our hold-out testing dataset. To enhance the performance of foundation models for segmenting kidney nuclei, we utilized multiple foundation models in a collective rating-based curation process and used the results as data enrichment strategies during fine-tuning. Our experimental results show that all three foundation models improved over their baselines, with StarDist achieving the highest F1 score of 0.82. Both StarDist and CellViT demonstrate that using “good” and “bad” image patches is the most effective training strategy. Notably, Cellpose, primarily a fluorescence-based model, raised its F1 score from 0.67 to 0.75 by incorporating “good” predictions from the other models. This mutual enhancement highlights the effectiveness of our strategy in improving segmentation outcomes. The improved model can further be applied in QuPath software to facilitate more efficient workflows in clinical kidney pathology.
Disclosures
The authors of the paper have no conflicts of interest to report.
Code, Data, and Materials Availability
Code will be coming soon and publicly available at https://github.com/hrlblab/AFM_kidney_cells.
Data will be provided after satisfying Vanderbilt University and Vanderbilt University Medical Center’s data use agreement.
Acknowledgments
This work was supported by the National Institutes of Health under award numbers R01EB017230, R01DK135597, T32EB001628, K01AG073584, and 5T32GM007347, and in part by the National Center for Research Resources and Grant UL1 RR024975-01. This study was also supported by the National Science Foundation (1452485, 1660816, and 1750213). The Vanderbilt Institute for Clinical and Translational Research (VICTR) is funded by the National Center for Advancing Translational Sciences (NCATS) Clinical Translational Science Award (CTSA) Program, Award Number 5UL1TR002243- 03. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH or NSF. This work was also supported by Vanderbilt Seed Success Grant, Vanderbilt Discovery Grant, and VISE Seed Grant. We extend gratitude to NVIDIA for their support by means of the NVIDIA hardware grant. This work was also supported by NSF NAIRR Pilot Award NAIRR240055.
The KPMP is funded by the following grants from the NIDDK: U01DK133081, U01DK133091, U01DK133092, U01DK133093, U01DK133095, U01DK133097, U01DK114866, U01DK114908, U01DK133090, U01DK133113, U01DK133766, U01DK133768, U01DK114907, U01DK114920, U01DK114923, U01DK114933, U24DK114886, UH3DK114926, UH3DK114861, UH3DK114915, UH3DK114937. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. The results here are in whole or part based upon data generated by the HuBMAP Program: https://hubmapconsortium.org. The Nephrotic Syndrome Study Network Consortium (NEPTUNE), U54-DK-083912, is a part of the National Institutes of Health (NIH) Rare Disease Clinical Research Network (RDCRN), supported through a collaboration between the Office of Rare Diseases Research (ORDR), NCATS, and the National Institute of Diabetes, Digestive, and Kidney Diseases. Additional funding and/or programmatic support for this project has also been provided by the University of Michigan, the NephCure Kidney International and the Halpin Foundation. The views expressed in written materials or publications do not necessarily reflect the official policies of the Department of Health and Human Services; nor does mention by trade names, commercial practices, or organizations imply endorsement by the U.S. Government.
Supplementary Information
Supplementary Figures S1, S2, and S3. Additional Qualitative Results of Enhanced Kidney Cell Nuclei Instance Segmentation. In addition to (A)-(C) shown in the Fig. 8, (D)-(L) are provided in Supplementary Figures S1 to S3.
Biographies and photographs of the other authors are not available.